OpenStudy (anonymous):

m-18=3m...solve for m

OpenStudy (anonymous):

-18=2m -9=m

OpenStudy (anonymous):

wht steps did yu take to do it?

OpenStudy (anonymous):

-18=3m-m -18=2m -18/2=m -9=m